OpenStudy (anonymous):

Please help!! Need help with five questions out of of 20 questions Find the value for tan(x/2) if the following conditions hold: cos(x) = 5/11 and x in in Quadrant IV A. - 2sqrt(6) / 3 B. sqrt(6)/4 C. - sqrt(6)/4 D. 2sqrt(6) / 3

OpenStudy (anonymous):

\[\cos x=\frac{ 1-\tan ^2\frac{ x }{ 2 } }{ 1+\tan ^2\frac{ x }{ 2 } }\]

OpenStudy (anonymous):

plug the value of cos x and cross multiply

OpenStudy (anonymous):

\[\tan \left( \frac{ x }{ 2 } \right)=\frac{ 1-\cos x }{ \sin x }\]

OpenStudy (anonymous):

you can find sin x from pythagorean theorem